OpenStudy (anonymous):

Will give medal and. Fan. Find the initial value problem of d^2y/dx ^2 + x dy/dx + y=0 given that y(0)=1 and dy/dx (0)=0

OpenStudy (anonymous):

Notice that \[\frac{d^2y}{dx^2}=\frac{d}{dx}\left[\frac{dy}{dx}\right]\quad\text{and}\quad x\frac{dy}{dx}+y=\frac{d}{dx}\left[xy\right]\] So, the original ODE is equivalent to \[\frac{d}{dx}\left[\frac{dy}{dx}+xy\right]=0\] Integrating both sides yields \[\frac{dy}{dx}+xy=C\] and you have an equation linear in \(y\). Find your integrating factor. \[\ln\mu(x)=\int x\,dx~~\implies~~\mu(x)=e^{x^2/2}\] Now solve, \[e^{x^2/2}\frac{dy}{dx}+xe^{x^2/2}y=Ce^{x^2/2}\]

OpenStudy (anonymous):

please can you help me finish it sir. it is an exam question . i dont want to do any error

OpenStudy (anonymous):

help me sir